# On computational complexity and average-case hardness of shallow-depth boson sampling

### v1

@@ V Average-case hardness implies classical simulation hardness
$||\mathcal{D}_{C}^{\prime}-\mathcal{D}_{C}||\leq\beta$ (8)
@@
We leave the proof of Lemma 4 in Appendix D for a more self-contained analysis. The complexity $\rm{BPP}^{\rm{NP}}$ is known to be inside the finite level of PH, i.e., $\rm{BPP}^{\rm{NP}}\subseteq PH$. Also, by Toda’s theorem [39], PH problems can be solved given the ability to solve any #P problem, i.e., $\rm{BPP}^{\rm{NP}}\subseteq\rm{PH}\subseteq\rm{P}^{\rm{\#P}}$. If finding the average-case solution of output probabilities of sampler $\mathcal{S}$ is #P-hard, then $\rm{P}^{\rm{\#P}}\subseteq\rm{BPP}^{\rm{NP}^{\mathcal{S}}}$. Therefore, if an efficient classical algorithm exists that can simulate $\mathcal{S}$, then $\rm{P}^{\rm{\#P}}\subseteq\rm{BPP}^{\rm{NP}}$ which implies the collapse of PH. Consequently, under the assumption of the non-collapse of the PH, there is no efficient classical algorithm capable of simulating $\mathcal{S}$.
Based on the above arguments, we show that for the case that allowed additive imprecision of Theorem 4 for the hardness can be improved, then it is classically hard to simulate shallow-depth boson sampling within a constant total variation distance.
@@
Suppose that the allowed additive imprecision for the problem in Theorem 4 to be #$\rm{P}$-hard can be improved to $\epsilon=2^{-(\gamma-1)N\log N-O(N)}$. Then the efficient classical simulation of approximate boson sampler $\mathcal{S}$ with respect to circuits from the shallow architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$ implies the collapse of $\rm{PH}$.
@@
For the randomly chosen circuit input, bad outcomes can vary with the circuit instances, as the sampler $\mathcal{S}$ has the freedom to choose its error distribution according to the input circuit. However, no matter how the bad outcomes vary with circuit instances, $\mathcal{O}$ succeeds at least $1-\frac{\xi}{\eta}$ fraction over circuit instances for at least $1-\eta$ fraction of the outcomes, for any $\eta$ satisfying $\xi<\eta<1$. Otherwise, the failure probability over outcomes and circuits would exceed $\xi$, which contradicts the proposition that the success probability of $\mathcal{O}$ is at least $1-\xi$ over randomly chosen outcomes and circuits. As the above argument holds for any random circuit families, we choose the random circuit distribution as $\mathcal{H}_{\mathcal{A}}$ with the shallow architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$.
@@
$\kappa\binom{M}{N}^{-1}$ $=\kappa\frac{N!(M-N)!}{M!}$ (10)
$=2^{-(\gamma-1)N\log N-O(N)},$ (11)
@@ VI Classical simulation hardness of shallow-depth Gaussian boson sampling
In this section, we show that our hardness results of the shallow-depth boson sampling can be generalized to the Gaussian boson sampling scheme  [2]. Our specific setup for the Gaussian boson sampling is as follows. Let the total mode number $M$ of the circuit be a power of 2, and now the input state is an $M$ product of single-mode squeezed vacuum (SMSV) state |$\ket{\text{SMSV}}^{\otimes M}$ with equal squeezing parameter $r$ and equal squeezing direction. Also, let us define the output mean photon number as an integer $N$ (i.e., $N=M\sinh^{2}r$) where $M$ and $N$ are polynomially related as $M=c_{1}N^{\gamma}$ for a constant $c_{1}$ and $\gamma\geq 1$. We define $q_{\bm{s}}(C)$ as an output probability of the Gaussian boson sampling, for an $N$ photon outcome $\bm{s}$ from an $M$ mode linear optical circuit matrix $C$ on input $M$ SMSV states. For collision-free outcome $\bm{s}$, $q_{\bm{s}}(C)$ can be expressed as [2]
@@
where |$\ket{\bm{s}}$ is an $M$-mode Fock state corresponding to the outcome $\bm{s}$, $\hat{\mathcal{U}}(C)$ is a unitary operator corresponding to the circuit $C$, and $(CC^{T})_{\bm{s}}$ is an $N$ by $N$ matrix obtained by taking $s_{i}$ copies of the $i$th row and column of the matrix $CC^{T}$.
@@
where the summation of $\bm{s}_{n}$ is over all possible configurations of Fock state with a total $M_{0}$ mode and $n$ photon.
For each mode in the given $M_{0}$ mode circuit $C_{0}$, one-half of the TMSV state (i.e., subscript (2) in Eq. (14)) is input into it, and the other half of each state (i.e., subscript (1) in Eq. (14)) is sent directly to a photon counter. By setting each |$\ket{\bm{s}_{\text{in}}}$ and |$\ket{\bm{s}_{\text{out}}}$ as a total $M_{0}$ mode and total $N_{0}$ photon Fock state, the output probability can be represented as
@@
Note that two $M_{0}$ mode $\mathcal{B}\mathcal{B}^{*}$ architecture can be embedded in the middle of an $M=2M_{0}$ mode $\mathcal{B}\mathcal{B}^{*}$ architecture. Accordingly, we define a circuit $C$ in $M$ mode $\mathcal{B}\mathcal{B}^{*}$ by embedding the given $M_{0}$ mode circuit $C_{0}$ in one side of $\mathcal{B}\mathcal{B}^{*}$, setting gates located right in front of the input ports as balanced beam splitters, and setting the remaining gates as identity gates; we leave in Fig. 3 an illustration of $M=16$ mode circuit $C$ for more clarity. Here, the input $M$ SMSV states with squeezing parameter $r$ combined with the balanced beam splitters at the front becomes $M_{0}$ TMSV states with squeezing parameter $r$. Therefore, our overall setup exactly follows the scheme in Ref. [40], such that the first $M_{0}$ mode is the photon counter sector to determine the input configuration of boson sampling, and the last $M_{0}$ mode is to simulate the boson sampling for the given circuit $C_{0}$.
@@
We also define an $M$-dimensional vector $\bm{s}$ as a serial concatenation of two $\bm{s}_{0}$ vectors, so that $\bm{s}$ represents $N$ photon outcome over $M$ modes. Then the output probabilities $p_{\bm{s}_{0}}(C_{0})$ and $q_{\bm{s}}(C)$ are related as
@@
Using the results of Theorem 6 and the previous proof of the average-case hardness of boson sampling in Theorem 4, it is straightforward to find the average-case hardness of Gaussian boson sampling, for randomly chosen $N$ photon outcomes $\bm{s}\sim\mathcal{G}_{M,N}$ and randomly chosen circuits $U\sim\mathcal{H}_{\mathcal{A}}$ in shallow-depth architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$.
@@
The following problem is #$\rm{P}$-hard under a $\rm{BPP}^{\rm{NP}}$ reduction: for any constant $\delta,\eta\geq 0$ with $\delta+\eta<\frac{1}{4}$, on input a random circuit $U\sim\mathcal{H}_{\mathcal{A}}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$ and a random outcome $\bm{s}\sim\mathcal{G}_{M,N}$, compute the output probability $q_{\bm{s}}(U)$ of Gaussian boson sampling within additive imprecision $\epsilon=2^{-O(N^{\gamma+1}(\log N)^{2})}$, with probability at least $1-\delta$ over the choice of $U$ for at least $1-\eta$ over the choice of $\bm{s}$.
Proof.
The procedure is the same as the proof of Theorem 4, namely, establishing a worst-to-average-case reduction from the problem in Theorem 6 to the problem in Theorem 7. The only different part for the Gaussian boson sampling case is the functional form of the output probability $q_{\bm{s}}(U(\theta))$ parameterized by $\theta$. Hence, we show that $q_{\bm{s}}(U(\theta))$ can also be represented as a degree $(4mN,4mN)$ rational function in $\theta$, the same degree as the boson sampling case in Lemma 3.
From Eq. (12), the output probability $q_{\bm{s}}(U(\theta))$ has the form of
@@
where $\mu$ is along all possible perfect matching permutations over $N$ modes. From the proof of Lemma 3, using reduction to the common denominator for all of the $m=qM\log M$ gates, $\left[U(\theta)\right]_{j,k}$ can be represented as $(2m,2m)$ rational function in $\theta$ with the common denominator $\prod_{i=1}^{m}q_{i}(\theta)$. Using this fact, one can easily check that $\prod_{j=1}^{N/2}\left[(U(\theta)U(\theta)^{T})_{\bm{s}}\right]_{\mu(2j-1),\mu(2j)}$ can be represented as $(2mN,2mN)$ rational function in $\theta$, with the common denominator $[\prod_{i=1}^{m}q_{i}(\theta)]^{N}$ which does not change with $\mu$. Therefore, the output probability can be represented as $q_{\bm{s}}(U(\theta))=\frac{P(\theta)}{Q(\theta)}$, with each $Q(\theta)=[\prod_{i=1}^{m}|q_{i}(\theta)|^{2}]^{N}$ and $P(\theta)$ a degree $4mN$ polynomial function in $\theta$.
Given that $q_{\bm{s}}(U(\theta))$ can be represented as a degree $(4mN,4mN)$ rational function with the same denominator $Q(\theta)=[\prod_{i=1}^{m}|q_{i}(\theta)|^{2}]^{N}$ from the boson sampling case in Lemma 3, we can repeat all the steps identically to the proof of Theorem 4 and obtain the same result. ∎
@@ VII Extension of hardness results for lossy environments
Given Eq. (20), we can repeat the same steps from the previous hardness arguments, for the lossy shallow-depth boson sampling; the only difference is the imprecision blowup by $2^{O(\rho m)}$. For our shallow-depth architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$, the gate number $m$ is $qM\log M$, so the size of imprecision blowup is $2^{O(N^{\gamma}\log N)}$ in our case. Such imprecision blowup does not affect the allowed additive accuracy $\epsilon=2^{-O(N^{\gamma+1}(\log N)^{2})}$ for our average-case hardness result. Based on the arguments so far, the following corollary is straightforward.
@@
Suppose we have the photon loss model $\mathcal{N}$ with each loss rate $\rho_{i}\leq\rho$ for a constant $\rho$. Then the following problem is #$\rm{P}$-hard under a $\rm{BPP}^{\rm{NP}}$ reduction: for any constant $\delta,\eta\geq 0$ with $\delta+\eta<\frac{1}{4}$, on input a random circuit $U\sim\mathcal{H}_{\mathcal{A}}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$ and a random outcome $\bm{s}\sim\mathcal{G}_{M,N}$, compute the lossy output probability $p_{\bm{s}}(U,\mathcal{N})$ within additive imprecision $\epsilon=2^{-O(N^{\gamma+1}(\log N)^{2})}$, with probability at least $1-\delta$ over the choice of $U$ for at least $1-\eta$ over the choice of $\bm{s}$.

### v2

@@ V Average-case hardness implies classical simulation hardness
$||\mathcal{D}_{C}^{\prime}-\mathcal{D}_{C}||\leq\beta,$ (3)
@@
We leave the proof of Lemma 2 in Appendix F for a more self-contained analysis. The complexity $\rm{BPP}^{\rm{NP}}$ is known to be inside the finite level of PH, i.e., $\rm{BPP}^{\rm{NP}}\subseteq PH$. Also, by Toda’s theorem [48], PH problems can be solved given the ability to solve any #P problem, i.e., $\rm{BPP}^{\rm{NP}}\subseteq\rm{PH}\subseteq\rm{P}^{\rm{\#P}}$. If finding the average-case solution of output probabilities of sampler $\mathcal{S}$ is #P-hard, then $\rm{P}^{\rm{\#P}}\subseteq\rm{BPP}^{\rm{NP}^{\mathcal{S}}}$. Therefore, if an efficient classical algorithm exists that can simulate $\mathcal{S}$, then $\rm{P}^{\rm{\#P}}\subseteq\rm{BPP}^{\rm{NP}}$ which implies the collapse of PH. Consequently, under the assumption of the non-collapse of the PH, there is no efficient classical algorithm capable of simulating $\mathcal{S}$.
Based on the above arguments, we show that for the case where the allowed additive imprecision of Theorem 4 for the hardness can be improved, then it is classically hard to simulate shallow-depth Boson Sampling within a constant total variation distance.
@@
Suppose that the allowed additive imprecision for the problem in Theorem 4 to be #$\rm{P}$-hard can be improved to $\epsilon=2^{-(\gamma-1)N\log N-O(N)}$. Then, in the dilute regime such that $\gamma\geq 2$, the efficient classical simulation of approximate boson sampler $\mathcal{S}$ with respect to circuits from the shallow architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 4}$ implies the collapse of $\rm{PH}$.
@@
Note that, Lemma 2 holds for a randomly chosen input circuit drawn from an arbitrary circuit ensemble, such that the success probability of $\mathcal{O}$ is at least $1-\xi$ over randomly chosen outcomes and circuits. However, for the randomly chosen circuit input, bad outcomes can vary with the circuit instances, as the sampler $\mathcal{S}$ has the freedom to choose its error distribution according to the input circuit (e.g., $C$ and $\bm{P}C$ would have different bad outcomes for permutation $\bm{P}$). Nevertheless, no matter how the bad outcomes vary with circuit instances, $\mathcal{O}$ succeeds at least $1-\frac{\xi}{\eta}$ fraction over circuit instances for at least $1-\eta$ fraction of the outcomes, for any $\eta$ satisfying $\xi<\eta<1$. Otherwise, the failure probability over outcomes and circuits would exceed $\xi$, which contradicts the proposition that the success probability of $\mathcal{O}$ is at least $1-\xi$ over randomly chosen outcomes and circuits. As the above argument holds for any random circuit families, we choose the random circuit distribution as $\mathcal{H}_{\mathcal{A}}\mathcal{P}$ with the shallow architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 3}$.
@@
$\kappa\binom{M}{N}^{-1}$ $=\kappa\frac{N!(M-N)!}{M!}=2^{-(\gamma-1)N\log N-O(N)},$ (5)
@@
We lastly remark that, because our average-case hardness result in Sec. IV solely relies on collision-free outcomes, it must be ensured that these collision-free outcomes occupy a sufficient portion of all possible outcomes of Boson Sampling (for more details, see Appendix B). Hence, we focus on the dilute regime $\gamma\geq 2$ for the classical hardness result in this section, to ensure the dominance of collision-free outcomes. To similarly obtain the classical hardness result in the saturated regime $21\leq\gamma<$ 2, extending our average-case hardness result for the collision outcomes is crucial.
@@ VI Classical simulation hardness of shallow-depth Gaussian Boson Sampling
In this section, we show that our hardness results of the shallow-depth Boson Sampling can be generalized to the Gaussian Boson Sampling scheme  [2]. Our specific setup for the Gaussian Boson Sampling is as follows. Let the total mode number $M$ of the circuit be a power of 2, and now the input state is an $M$ product of single-mode squeezed vacuum (SMSV) state |$\ket{\text{SMSV}}^{\otimes M}$ with equal squeezing parameter $r$ and equal squeezing direction. Also, let us define the output mean photon number as an integer $N$ (i.e., $N=M\sinh^{2}r$) where $M$ and $N$ are polynomially related as $M=c_{1}N^{\gamma}$ for a constant $c_{1}$ and $\gamma\geq 2$. We define $q_{\bm{s}}(C)$ as an output probability of the Gaussian Boson Sampling, for an $N$-photon outcome $\bm{s}$ from an $M$-mode linear optical circuit matrix $C$ on input $M$ SMSV states. For collision-free outcome $\bm{s}$, $q_{\bm{s}}(C)$ can be expressed as [2]
@@
where |$\ket{\bm{s}}$ is an $M$-mode Fock-state corresponding to the outcome $\bm{s}$, $\hat{\mathcal{U}}(C)$ is a unitary operator corresponding to the circuit $C$, and $(CC^{T})_{\bm{s}}$ is an $N$ by $N$ matrix obtained by taking $s_{i}$ copies of the $i$th row and column of the matrix $CC^{T}$.
@@
where the summation of $\bm{s}_{n}$ is over all possible configurations of Fock-state with a total $M_{0}$ mode and $n$ photon.
For each mode in the given $M_{0}$ mode circuit $C_{0}$, one-half of the TMSV state (i.e., subscript (2) in Eq. (8)) is input into it, and the other half of each state (i.e., subscript (1) in Eq. (8)) is sent directly to a photon counter. By setting each |$\ket{\bm{s}_{\text{in}}}$ and |$\ket{\bm{s}_{\text{out}}}$ as a total $M_{0}$ mode and total $N_{0}$ photon Fock-state, the output probability can be represented as
@@
Note that two $M_{0}$ mode $\mathcal{B}\mathcal{B}^{*}$ architecture can be embedded in the middle of an $M=2M_{0}$ mode $\mathcal{B}\mathcal{B}^{*}$ architecture. Accordingly, we define a circuit $C$ in an $M$-mode $\mathcal{B}\mathcal{B}^{*}$ by embedding the given $M_{0}$ mode circuit $C_{0}$ in one side of $\mathcal{B}\mathcal{B}^{*}$, setting gates located right in front of the input ports as balanced beam splitters, and setting the remaining gates as identity gates; we leave in Fig. 3 an illustration of $M=16$ mode circuit $C$ for more clarity. Here, the input $M$ SMSV states with squeezing parameter $r$ combined with the balanced beam splitters at the front becomes $M_{0}$ TMSV states with squeezing parameter $r$. Therefore, our overall setup exactly follows the scheme in Ref. [49], such that the first $M_{0}$ mode is the photon counter sector to determine the input configuration of Boson Sampling, and the last $M_{0}$ mode is to simulate the Boson Sampling for the given circuit $C_{0}$.
@@
We also define an $M$-dimensional vector $\bm{s}$ as a serial concatenation of two $\bm{s}_{0}$ vectors, so that $\bm{s}$ represents an $N$-photon outcome for $M$ modes. Then the output probabilities $p_{\bm{s}_{0}}(C_{0})$ and $q_{\bm{s}}(C)$ are related as
@@
Using the results of Theorem 6 and the previous proof of the average-case hardness of Boson Sampling in Theorem 4, it is straightforward to find the average-case hardness of Gaussian Boson Sampling, for a randomly chosen $N$-photon outcomes $\bm{s}\sim\mathcal{G}_{M,N}$ and a randomly chosen circuit $U\sim\mathcal{H}_{\mathcal{A}}$ in shallow-depth architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$.
@@
The following problem is #$\rm{P}$-hard under a $\rm{BPP}^{\rm{NP}}$ reduction: for any constant $\delta,\eta\geq 0$ with $\delta+\eta<\frac{1}{4}$, on input a $O(\log N)$-depth random circuit $U\sim\mathcal{H}_{\mathcal{A}}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$ and a random collision-free outcome $\bm{s}\sim\mathcal{G}_{M,N}$, compute the output probability $q_{\bm{s}}(U)$ of Gaussian Boson Sampling within additive imprecision $\epsilon=2^{-O(N^{\gamma+1}(\log N)^{2})}$, with probability at least $1-\delta$ over the choice of $U$ for at least $1-\eta$ over the choice of $\bm{s}$.
It is worth emphasizing that for our setup of the Gaussian Boson Sampling, the bosonic birthday paradox is already guaranteed in the dilute regime $M=\Omega(N^{2})$, in contrast to the Boson Sampling case. More specifically, the bosonic birthday paradox requires the symmetry of inputs over the unitary circuits, such that the input pattern should be invariant under the permutations (see Appendix B for more details). Here, note that because we are considering “full mode” input, our setup already satisfies this required input symmetry. To see this symmetry clearly, one can easily check that for any unitary circuit $U$ and permutation matrix $\bm{P}$, the output probability of Gaussian Boson Sampling for $U\bm{P}$ is equivalent to
$q_{\bm{s}}(U\bm{P})=\frac{\tanh^{N}r}{\cosh^{M}r}|\text{Haf}((U\bm{P}\bm{P}^{T}U^{T})_{\bm{s}})|^{2}=\frac{\tanh^{N}r}{\cosh^{M}r}|\text{Haf}((UU^{T})_{\bm{s}})|^{2}=q_{\bm{s}}(U).$ (12)
Hence, we do not necessarily have to include the random permutation part $\mathcal{P}$ for the average-case hardness result; only the local random circuit for $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$ suffices.
Proof of Theorem 7.
The procedure is essentially the same as the proof of Theorem 4, namely, establishing a worst-to-average-case reduction from the problem in Theorem 6 to the problem in Theorem 7: First, given worst-case circuit $C\in(\mathcal{B}\mathcal{B})^{q\geq 1}$ and outcome $\bm{s}_{0}$ depicted in Theorem 6, sample permutation $\bm{P}$, such that the revised worst-case circuit and outcome become $C_{p}=\bm{P}C\in(\mathcal{B}\mathcal{B})^{q\geq 2}$ and $\bm{s}=\bm{P}\bm{s}_{0}\sim\mathcal{G}_{M,N}$ respectively. Then, we sample $V(\theta)$ from local random circuit ensemble with circuit architecture $\mathcal{A}=(\mathcal{B}\mathcal{B})^{q\geq 2}$ and perturb by $C_{p}=\bm{P}C$ via Cayley transform, such that $V(\theta)$ follows the distribution $\mathcal{H}_{\mathcal{A},\theta}^{C_{p}}$. Given access to the oracle that solves the problem in Theorem 7, we infer the output probability $q_{\bm{s}}(V(\theta))$ using different values of $\theta$, and thus obtain the worst-case output probability $q_{\bm{s}}(V(1))=q_{\bm{s}}(C_{p})=q_{\bm{s}_{0}}(C)$ (for more details, see Appendix E, where we prove the average-case hardness for the Boson Sampling case).
Hence, the only different part for the Gaussian Boson Sampling case is the functional form of the output probability $q_{\bm{s}}(V(\theta))$ parameterized by $\theta$ via Cayley transformation. Accordingly, we show that $q_{\bm{s}}(V(\theta))$ can also be represented as a degree $(4mN,4mN)$ rational function in $\theta$, the same degree as the Boson Sampling case in Lemma 9.
From Eq. (6), the output probability $q_{\bm{s}}(V(\theta))$ has the form of
@@
where $\mu$ is along all possible perfect matching permutations over $N$ modes. From the proof of Lemma 9, using reduction to the common denominator for all of the $m=qM\log M$ gates, $\left[V(\theta)\right]_{j,k}$ can be represented as $(2m,2m)$ rational function in $\theta$ with the common denominator $\prod_{i=1}^{m}q_{i}(\theta)$. Using this fact, one can easily check that $\prod_{j=1}^{N/2}\left[(V(\theta)V(\theta)^{T})_{\bm{s}}\right]_{\mu(2j-1),\mu(2j)}$ can be represented as $(2mN,2mN)$ rational function in $\theta$, with the common denominator $[\prod_{i=1}^{m}q_{i}(\theta)]^{N}$ which does not change with $\mu$. Therefore, the output probability can be represented as $q_{\bm{s}}(V(\theta))=\frac{P(\theta)}{Q(\theta)}$, with each $Q(\theta)=[\prod_{i=1}^{m}|q_{i}(\theta)|^{2}]^{N}$ and $P(\theta)$ a degree $4mN$ polynomial function in $\theta$.
Given that $q_{\bm{s}}(V(\theta))$ can be represented as a degree $(4mN,4mN)$ rational function with the same denominator $Q(\theta)=[\prod_{i=1}^{m}|q_{i}(\theta)|^{2}]^{N}$ from the Boson Sampling case in Lemma 9, we can repeat all the steps identically to the proof of Theorem 4 and obtain the same result. ∎
@@ VII Extension of hardness results for lossy environments
Given Eq. (15), we can repeat the same steps from the previous hardness arguments, for the lossy shallow-depth Boson Sampling; the only difference is the imprecision blowup by $2^{O(\rho m)}$. For our shallow-depth architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$, the gate number $m$ is $qM\log M$, so the size of imprecision blowup is $2^{O(N^{\gamma}\log N)}$ in our case. Such imprecision blowup does not affect the allowed additive accuracy $\epsilon=2^{-O(N^{\gamma+1}(\log N)^{2})}$ for our average-case hardness result. Based on the arguments so far, the following corollary is straightforward from Theorem 4.
@@
Suppose we have the photon loss model $\mathcal{N}$ with each loss rate $\rho_{i}\leq\rho$ for a constant $\rho$. Then the following problem is #$\rm{P}$-hard under a $\rm{BPP}^{\rm{NP}}$ reduction: for any constant $\delta,\eta\geq 0$ with $\delta+\eta<\frac{1}{4}$, on input a $O(\log N)$-depth random circuit $U\sim\mathcal{H}_{\mathcal{A}}\mathcal{P}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 3}$ and a random collision-free outcome $\bm{s}\sim\mathcal{G}_{M,N}$, compute the lossy output probability $p_{\bm{s}}(U,\mathcal{N})$ within additive imprecision $\epsilon=2^{-O(N^{\gamma+1}(\log N)^{2})}$, with probability at least $1-\delta$ over the choice of $U$ for at least $1-\eta$ over the choice of $\bm{s}$.